# Solving Nonlinear PDEs with Sparse Radial Basis Function Networks

### v1

@@ 4.1.3. Boundary conditions treatments
Figure 3. Numerical results of different treatments of boundary conditions. (A) Exact solution $u(x_{1},x_{2})$; (B) Error and convergence speed – penalty coefficient $\lambda$ (results are averaged from 10 runs); (C) Error contour of solutions with $L^{2}$-type loss, $H^{1}$-type loss, and “mask” function for boundary conditions. Each $\star$ represents the location parameters $y_{n}$ of learned kernel nodes, and the dotted circle around are of radius $\sigma_{n}$. The contour plots are taken from a single run, while the reported number of iterations, $L^{2}$ error and $L^{\infty}$ error are averaged over 10 runs.
@@ 4.2. Eikonal equation
Table 4. Errors and residual loss of numerical solutions to the regularized Eikonal equation ($\epsilon=0.10$) using grid collocation points with spatial mask $\gamma=(1-x_{1}^{2})(1-x_{2}^{2})$. Test results are evaluated on a $100\times 100$ grid in $D$. All results are averaged over 10 runs.
@@
We then check the convergence towards the unique viscosity solution as $\epsilon\rightarrow 0$. The unique viscosity solution in this case is $u_{\text{visc}}=\text{dist}(x,\partial D)=\min(1-|x_{1}|,1-|x_{2}|)$. We solve the Eikonal equation with $\epsilon=0.5,0.1,0.05,0.01$ respectively with $K=30^{2}$ grid collocation points with mask function $\gamma=(1-x_{1}^{2})(1-x_{2}^{2})$ and $\alpha=10^{-6}$. We observe that our method shows excellent convergence behavior as $\epsilon$ approaches 00 (see Figure 5). As the solution becomes increasingly non-smooth near the diagonals, kernel nodes adaptively cluster in these regions, capturing the sharp transition of the solution.
Figure 4. Numerical results of regularized Eikonal equation ($\epsilon=0.10$). (A) Exact solution $u(x_{1},x_{2})$, $\epsilon=0.10$; (B) Error contour of numerical solutions with regularization parameter $\alpha=10^{-4},10^{-6}$. Each $\star$ represents the location parameters $y_{n}$ of learned kernel nodes, and the dotted circle around are of radius $\sigma_{n}$.
Figure 5. Convergence of numerical solution to viscosity solution $u_{\text{visc}}(x)=\text{dist}(x,\partial D)=1-\max(|x_{1}|,|x_{2}|)$ as $\epsilon\rightarrow 0$. (A) Solutiuon surfaces (shadow in the background represents viscosity solution); (B) Error contours with shrinking $\epsilon$. Each $\star$ represents the location parameters $y_{n}$ of learned kernel nodes, and the dotted circle around are of radius $\sigma_{n}$. The contour plots are taken from a single run, while the reported $L^{2}$ and $L^{\infty}$ errors are averaged over 10 runs.
4.3. Viscous Burgers’ equation
@@
Instead of using a spatial-temporal formulation as in [15, 47], we employ a simple (fully) implicit backward Euler method for time discretization, which not only reduces the dimension of the problem but also alleviates the difficulty of selecting appropriate penalty parameter $\lambda$ that simultaneously balances initial and boundary conditions. The resulting algorithm solves an array of one-dimensional equations for $\{u^{n}\}_{n=1}^{N}$ ($N=\lceil\frac{1}{\Delta t}\rceil$) sequentially starting from $u^{0}=u(0,\cdot)$. To specify, we use Algorithm 1 at each time step $n$ with the residuals
@@
Figure 6. Viscous Burgers equation solved by Sparse RBFNet coupled with full-implict time discretization. (A) Convergence of loss for selected $u^{n}$ ($\alpha=10^{-4}$); (B) Error contour ($\alpha=10^{-4}$); (C) Growth of number of kernels in the network during time steps ($\Delta t=0.01$) for $\alpha=10^{-3},10^{-4}$ (results averaged from 10 runs); (D) Time slices of exact and true solution at time $t=0.15,0.50,0.85$ ($\alpha=10^{-4}$). Size of $\star$ is proportional to the bandwidth of individual kernel.
@@
Table 5. Errors and number of kernels for the viscous Burgers’ equation with different combinations of time step $\Delta t$ and regularization parameter $\alpha$. Numerical solution $\{u^{n}\}$ is extended to the entire $D$ by linear interpolation. Test results are evaluated on a grid of $120\times 120$ in $D$. The empirical residual loss $\hat{L}$ is not accessible because some test points do not lie on the discretized time steps. The number of kernels shown in the table is first averaged between time steps. All results are averaged over 10 runs.
@@
We note that refined time discretization schemes (e.g., higher-order backward differentiation, adaptive step sizing) can also be incorporated to further improve accuracy [16]. While valid, these points are out of the scope of this work and will be left as future research directions.
5. Conclusion and discussion
@@
Further improvement of the computational algorithm can be achieved as well through the usage of kernels and loss functions that better capture features and regularity of both the solution and source function data. For instance, the effectiveness of a well-chosen loss function is exemplified in Section 4.1.3 in which $H^{1}(\partial D)$ is used in place of the usual $L^{2}(\partial D)$ for boundary constraints. Additionally, as the computational cost of our method scales with the number of collocation points $K$, number of kernel functions $N$, and problem dimension $d$, it is natural to consider a doubly adaptive sub-sampling framework, where only a subset of the kernel parameters is updated with only a mini-batch of the collocation points at each iteration. While the use of mini-batches of training data is standard in the machine learning literature, incorporating it with partial kernel weights update (usually known as block coordinate descent method [63, 64]) requires further theoretical justification and algorithmic design. Nevertheless, this strategy can significantly improve the computational efficiency of the algorithm, particularly in cases where large $K$ and $N$ are required to capture complex solution geometries, which is common in high-dimensional PDE problems. Finally, it is also of great interest to explore extensions of this approach to inverse problems and nonlocal or fractional equations [12, 19, 29].
@@ A.1.1. Insertion threshold
A.1.2. Loosening insertion criterion
@@
(A.1) $\text{Pr}(\omega_{N+1})=\min\{1,\;\exp(-\frac{1}{T\cdot\hat{R}[u]}(\max_{1\leq n\leq N}|p[u](\omega_{n})|-|p[u](\omega_{N+1})|))\}$
@@
A.2. Phase II
A.2.1. Levenberg–Marquardt type correction term for invertability
@@ A.4. Pretraining
Figure 7. Convergence of residual loss for $\alpha=10^{-3}$ with/without pretraining.
@@ Appendix B Choice of collocation points
Table 6. Errors, residual loss, and number of kernels for Sparse RBFNet and GP using randomly generated collocation points (in comparison to Table 6). Case with $\alpha=10^{-3}$($10^{-4}$) is solved with solution obtained by $\alpha=10^{-2}$($10^{-3}$) as initialization. Test results are evaluated on a $100\times 100$ grid in $D$. All results are averaged over 10 runs. For the Gaussian Process method, we retain the data even if some runs diverge due to random sampling.
@@ Appendix C Besov space characterization
Theorem C.1.
Consider the unrestricted dictionary $\Omega=\mathbb{R}^{d}\times(0,\sigma_{\max}]$ with $\sigma_{\max}=1$. Then, the corresponding space
@@
admits a continuous embedding of the Besov space $B^{s}_{1,1}(\mathbb{R}^{d})$ (as defined in, e.g., [51, Section 2.1.1]) into $\mathcal{V}$.
@@
see [51, Section 2.1.1]. Here, $\psi\colon\mathbb{R}^{d}\to[0,1]$ is a function in $C_{c}^{\infty}(\mathbb{R}^{d})$ with
@@
(C.1) $u_{j}=\hat{\varphi}_{j}*\mu_{j}\quad\text{with }\mu_{j}\in L^{1}(\mathbb{R}^{d}).$
@@
$\lVert u\rVert_{\mathcal{V}(\mathbb{R}^{d})}\leq\lVert\mu\rVert_{M(\Omega)}=\sum_{j=0}^{\infty}2^{j\,s}\lVert\mu_{j}\rVert_{L^{1}}\leq c\sum_{j=0}^{\infty}2^{j\,s}\lVert u_{j}\rVert_{L^{1}}=c\lVert u\rVert_{B^{s}_{1,1}(\mathbb{R}^{d})}$

### v2

@@ 4.1.3 Boundary conditions treatments
Figure 3: Numerical results of different treatments of boundary conditions. (a) Exact solution $u(x_{1},x_{2})$; (b) Error and convergence speed – penalty coefficient $\lambda$ (results are averaged from 10 runs); (c) Error contour of solutions with $L^{2}$-type loss, $H^{1}$-type loss, and “mask” function for boundary conditions. Each $\star$ represents the location parameter $y_{n}$ of a learned kernel node, and the dotted circle around has radius $\sigma_{n}$. The contour plots are taken from a single run, while the reported number of iterations, $L^{2}$ error and $L^{\infty}$ error are averaged over 10 runs.
@@ 4.2 Eikonal equation
Table 4: Errors and residual loss of numerical solutions to the regularized Eikonal equation ($\epsilon=0.10$) using grid collocation points with spatial mask $\gamma(x)=(1-x_{1}^{2})(1-x_{2}^{2})$. Test results are evaluated on a $100\times 100$ grid in $D$. All results are averaged over 10 runs.
@@
We then check the convergence towards the unique viscosity solution as $\epsilon\rightarrow 0$. The unique viscosity solution in this case is $u_{\text{visc}}=\text{dist}(x,\partial D)=\min(1-|x_{1}|,1-|x_{2}|)$. We solve the Eikonal equation with $\epsilon=0.5,0.1,0.05,0.01$ respectively with $K=30^{2}$ grid collocation points with mask function $\gamma(x)=(1-x_{1}^{2})(1-x_{2}^{2})$ and $\alpha=10^{-6}$. We observe that our method shows excellent convergence behavior as $\epsilon$ approaches 00 (see Figure 5). As the solution becomes increasingly non-smooth near the diagonals, kernel nodes adaptively cluster in these regions, capturing the sharp transition of the solution.
Figure 4: Numerical results of regularized Eikonal equation ($\epsilon=0.10$). (a) Exact solution $u(x_{1},x_{2})$, $\epsilon=0.10$; (b) Error contour of numerical solutions with regularization parameter $\alpha=10^{-4},10^{-6}$. Each $\star$ represents the location parameters $y_{n}$ of learned kernel nodes, and the dotted circle around are of radius $\sigma_{n}$.
Figure 5: Convergence of numerical solution to viscosity solution $u_{\text{visc}}(x)=\text{dist}(x,\partial D)=1-\max(|x_{1}|,|x_{2}|)$ as $\epsilon\rightarrow 0$. (a) Solutiuon surfaces (shadow in the background represents viscosity solution); (b) Error contours with shrinking $\epsilon$. Each $\star$ represents the location parameters $y_{n}$ of learned kernel nodes, and the dotted circle around are of radius $\sigma_{n}$. The contour plots are taken from a single run, while the reported $L^{2}$ and $L^{\infty}$ errors are averaged over 10 runs.
4.3 Viscous Burgers’ equation
@@
Instead of using a spatial-temporal formulation as in Chen et al. (2021); Raissi et al. (2019), we employ a simple (fully) implicit backward Euler method for time discretization, which not only reduces the dimension of the problem but also alleviates the difficulty of selecting appropriate penalty parameter $\lambda$ that simultaneously balances initial and boundary conditions. The resulting algorithm solves an array of one-dimensional equations for $\{u^{n}\}_{n=1}^{N}$ ($N=\lceil\frac{1}{\Delta t}\rceil$) sequentially starting from $u^{0}=u(0,\cdot)$. To specify, we use Algorithm 1 at each time step $n$ with the residuals
@@
Figure 6: Viscous Burgers equation solved by Sparse RBFNet coupled with fully implicit time discretization. (a) Convergence of loss for selected $u^{n}$ ($\alpha=10^{-4}$); (b) Error contour ($\alpha=10^{-4}$); (c) Growth of number of kernels in the network during time steps ($\Delta t=0.01$) for $\alpha=10^{-3},10^{-4}$ (results averaged from 10 runs); (d) Time slices of exact and true solution at time $t=0.15,0.50,0.85$ ($\alpha=10^{-4}$). The size of $\star$ is proportional to the bandwidth of an individual kernel.
@@
Table 5: Errors and number of kernels for the viscous Burgers’ equation with different combinations of time step $\Delta t$ and regularization parameter $\alpha$. Numerical solution $\{u^{n}\}$ is extended to the entire $D$ by linear interpolation. Test results are evaluated on a grid of $120\times 120$ in $D$. The empirical residual loss $\hat{L}$ is not accessible because some test points do not lie on the discretized time steps. The number of kernels shown in the table is first averaged between time steps. All results are averaged over 10 runs.
@@
We note that refined time discretization schemes (e.g., higher-order backward differentiation, adaptive step sizing) can also be incorporated to further improve accuracy (Chen et al., 2025). While valid, these points are out of the scope of this work and will be left as future research directions.
5 Conclusion and discussion
@@
Further improvement of the computational algorithm can be achieved as well through the usage of kernels and loss functions that better capture features and regularity of both the solution and source function data. For instance, the effectiveness of a well-chosen loss function is exemplified in Section 4.1.3 in which $H^{1}(\partial D)$ is used in place of the usual $L^{2}(\partial D)$ for boundary constraints. Additionally, as the computational cost of our method scales with the number of collocation points $K$, number of kernel functions $N$, and problem dimension $d$, it is natural to consider a doubly adaptive sub-sampling framework, where only a subset of the kernel parameters is updated with only a mini-batch of the collocation points at each iteration. While the use of mini-batches of training data is standard in the machine learning literature, incorporating it with partial kernel weights update (usually known as block coordinate descent method (Wright, 2015; Xu and Yin, 2013)) requires further theoretical justification and algorithmic design. Nevertheless, this strategy can significantly improve the computational efficiency of the algorithm, particularly in cases where large $K$ and $N$ are required to capture complex solution geometries, which is common in high-dimensional PDE problems. Finally, it is also of great interest to explore extensions of this approach to inverse problems and nonlocal or fractional equations (Burkardt et al., 2021; D’Elia et al., 2020; Guo et al., 2022).
Acknowledgments and Disclosure of Funding
@@ A.1.1 Insertion threshold
A.1.2 Loosening insertion criterion
@@
$\text{Pr}(\omega_{N+1})=\min\{1,\;\exp(-\frac{1}{T\cdot\hat{R}[u]}(\max_{1\leq n\leq N}|p[u](\omega_{n})|-|p[u](\omega_{N+1})|))\}$ (A.1)
@@
A.2 Phase II
A.2.1 Levenberg–Marquardt type correction term for invertability
@@ A.4 Pretraining
Figure 7: Convergence of residual loss for $\alpha=10^{-3}$ with/without pretraining.
@@ Appendix B Choice of collocation points
Table 6: Errors, residual loss, and number of kernels for Sparse RBFNet and GP using randomly generated collocation points (in comparison to Table 6). Case with $\alpha=10^{-3}$($10^{-4}$) is solved with solution obtained by $\alpha=10^{-2}$($10^{-3}$) as initialization. Test results are evaluated on a $100\times 100$ grid in $D$. All results are averaged over 10 runs. For the Gaussian Process method, we retain the data even if some runs diverge due to random sampling.
@@ Appendix C Besov space characterization
Theorem 24.
Consider the unrestricted dictionary $\Omega=\mathbb{R}^{d}\times(0,\sigma_{\max}]$ with $\sigma_{\max}=1$. Then, for the corresponding space
@@
there is a continuous embedding of the Besov space $B^{s}_{1,1}(\mathbb{R}^{d})$ (as defined in, e.g., Sawano, 2018, Section 2.1.1) into $\mathcal{V}$.
@@
(see Sawano, 2018, Section 2.1.1). Here, $\psi\colon\mathbb{R}^{d}\to[0,1]$ is a function in $C_{c}^{\infty}(\mathbb{R}^{d})$ with
@@
$u_{j}=\hat{\varphi}_{j}*\mu_{j}\quad\text{with }\mu_{j}\in L^{1}(\mathbb{R}^{d}).$ (C.1)
@@
$\lVert u\rVert_{\mathcal{V}(\mathbb{R}^{d})}\leq\lVert\mu\rVert_{M(\Omega)}=\sum_{j=0}^{\infty}2^{j\,s}\lVert\mu_{j}\rVert_{L^{1}}\leq c\sum_{j=0}^{\infty}2^{j\,s}\lVert u_{j}\rVert_{L^{1}}=c\lVert u\rVert_{B^{s}_{1,1}(\mathbb{R}^{d})}.\qed$